A point charge q1 = -2.4 ?C is located at the origin of a co-ordinate system. Another point charge q2 = 6.1 ?C is located along the x-axis at a distance x2 = 9.8 cm from q1 and displaced a distance y2 = 2.1 cm in the positive y-direction.

A third point charge q3 is now positioned halfway between q1 and q2. The net force on q2 now has a magnitude of F2,net = 4.372 N and points away from q1 and q3. What is the value (sign and magnitude) of the charge q3?
Answer #1

Since the net force is away from q1 and q3, q3 should be +ve charge since force due to q1 points towards q1 and force due to q3 should point away from q3 so that the net force points away from q1 and q3.

F2,net = 4.372

d12 = sqrt(2.1^2 + 9.8^2) = 10.02 cm

d23 = d12/2 = 5.01 cm

F2,net = kq1q2/d12^2 + kq2q3/d23^2

4.372 = (9x10^9)x10^-12[-2.4*6.1/(0.1002)^2 + 6.1*q3/(0.0501)^2]

So q3 = +0.8 x 10^-6 C = +0.8 uC